Two planes leave LA at the same time. One heads south to San Diego, the other heads north to San Francisco. The San Francisco plane flies 50 mph faster. In 1/2 hr, the planes are 275 mi apart. That are their speeds?
Answer by Paul(988)

San F. plane = x+50
La plane = x
1/2hours = 0.5h
0.5(x+x+50)=275
x+25=275
x=250

250+50=300

Hence, the speed of one plane is 250mph and another is 300mph.
Paul.